if i make a 67 one six weeks and a 54 on the other what do i have to make in the third six weeks to avarage out at 70?
Found 2 solutions by Alan3354, unlockmath:
Answer by Alan3354(69443)

if i make a 67 one six weeks and a 54 on the other what do i have to make in the third six weeks to avarage out at 70?
-------
(67 + 54 + grade)/3 = 70
121 + grade = 210
grade = 89

Answer by unlockmath(1688)

Hello,
A bit confusing about the "3rd six weeks" but I think I get your drift. Try this:
(67+54+x)/3=70
Multiply by 3 and add to get:
121+x=210
Subtract 121:
x=89 dollars
RJ
www.math-unlock.com